# Energy-Based Model for Accurate Estimation of Shapley Values in Feature Attribution

### v3

@@ Abstract
Shapley value is a widely used tool in explainable artificial intelligence (XAI), as it provides a principled way to attribute contributions of input features to model outputs. However, estimation of Shapley value requires capturing conditional dependencies among all feature combinations, which poses significant challenges in complex data environments. In this article, EmSHAP (Energy-based model for Shapley value estimation), an accurate and efficient Shapley value estimation method, is proposed to estimate the expectation of Shapley contribution function under the arbitrary subset of features given the rest. By utilizing the ability of energy-based model (EBM) to model complex distributions, EmSHAP provides an effective solution for estimating the required conditional probabilities. To further improve estimation accuracy, a GRU (Gated Recurrent Unit)-coupled partition function estimation method is introduced. The GRU network captures long-term dependencies with a lightweight parameterization and maps input features into a latent space to mitigate the influence of feature ordering. Additionally, a dynamic masking mechanism is incorporated to further enhance the robustness and accuracy by progressively increasing the masking rate. Theoretical analysis on the error bound as well as application to four case studies verified the higher accuracy and better scalability of EmSHAP in contrast to competitive methods.
Index Terms:
Explainable artificial intelligence, feature attribution, Shapley value, energy-based model.
@@ I Introduction
Deep learning (DL) has been extensively applied in many complex supervised learning problems like medical diagnosis [1], industrial intelligence [2], and financial modeling [3]. DL models outperform conventional methods in terms of accuracy in many cases. However, since most DL models are black boxes and have complicated structures, interpretation of the results becomes difficult. This limitation poses challenges in identifying key features that influence the DL model and understanding the relationship between input features and outputs. This interpretability requirement leads to the development of an active research area of Explainable Artificial Intelligence (XAI) [4].
@@
A notable problem for Shapley-based explanation is its high computation load. By definition, the computational complexity of Shapley value increases exponentially with the number of variables. In order to accelerate computation, a series of statistical-based methods have been developed that can be roughly divided into three categories: regression-based methods, sampling-based methods, and structure model-based methods. Regression-based methods employ weighted linear regression to accelerate the estimation of Shapley value, with examples including SHAP (SHapley Additive exPlanation) [11] and KernelSHAP [12]. Sampling-based methods utilize Monte Carlo sampling to approximate the true Shapley value. These methods involve Monte Carlo approximation using randomly sampled permutations [13] or sampling based on random permutations of local features [14]. On the other hand, structure model-based methods leverage specific model structure constraints to approximate the Shapley value. Examples of such methods include tree Shapley [15] and graph Shapley [16]. More recently, a generative model for Shapley value estimation based on variational autoencoder with arbitrary conditioning (VAEAC) [17] has been developed, which uses a single model for all feature combinations.
@@
Despite the advancements made in calculation of Shapley value, several significant issues still remain. For example, regression-based methods can improve the estimation efficiency of Shapley value, however, the inherent assumption of independence among variables may negatively affect its accuracy. Sampling-based methods require a large quantity of samplings to achieve sufficient accuracy, and hence have high computation load. Whilst structure model-based approaches demonstrate improved efficiency, their application scenarios are limited to specific models. For VAEAC, it involves an additional amortization error by using the VAE model [18, 19]. Another notable research gap is that there still lacks rigid theoretical analysis on the error bounds of different Shapley value-based approaches, making it difficult to quantitatively evaluate their accuracies.
@@
In order to increase estimation accuracy and provide theoretical guidance on the error bound, this article presents the EmSHAP (Energy-based model SHapley value estimation). Similar to Ref. [17], EmSHAP introduces a generative model during the estimation of Shapley contribution function. It relies on the energy-based model(EBM) to obtain the expectation of Shapley contribution function of the deep learning model under some arbitrary subset of features given the rest. The use of EBM is motivated by its ability to model complex, high-dimensional distributions without explicit normalization, which provides a flexible and efficient framework for estimating the conditional distributions required in Shapley value computation. Unlike original EBM, the EBM in EmSHAP is decomposed into an energy network to approximate the unnormalized conditional density and a GRU network for approximation of proposal distribution in the partition function. The GRU network has a simple structure and is capable of capturing long-term dependencies, and the adopted dynamic masking scheme improves the flexibility. The contributions of the paper can be summarized: i) Methodological refinements on the EBM framework by introducing the GRU network and the dynamic masking scheme, resulting in more accurate Shapley value estimation; ii) Theoretical contribution on the analysis of error bounds for several state-of-the-art methods, which confirmed the tighter error bound of EmSHAP; iii) Extensive application experiments are conducted to verify the accuracy and scalability of different Shapley value estimation methods.
II Relate work
@@
Shapley value is a concept originated from cooperative game theory and provides a fair and reasonable method for quantifying feature attribution contributions. As a result, they have garnered significant attention in the field of XAI. However, the calculation of exact Shapley values is an NP-hard problem [20]. To address this computational complexity, researchers have developed various approximation methods. This section provides a brief review on the different kinds of approximation methods in literature, which can be summarized into sampling-based, regression-based, gradient-based, and generative model-based methods.
@@
Sampling-based methods: Sampling-based Shapley value estimation is among the first to be studied and remains the most widely used. The core principle is to randomly sample subsets of features and calculate the average contribution of the features to model output. Variants of sampling-based methods include Monte Carlo sampling [21], Monte Carlo sampling with random arrangements [22, 7], and Owen-based sampling [23]. Most of them are model-agnostic and do not involve feature removal in Shapley value estimation, hence are popular in practical applications. However, sampling-based methods may suffer from high variance during the sampling procedures. The recently proposed SVARM [24] addresses this issue by controlling the estimated variance within a reasonable range.
@@
Regression-based methods: Regression-based methods estimate Shapley values by using regression models to approximate the marginal contribution of each feature to the output. These methods approximate the original black-box model by constructing a simpler surrogate model. Among these methods, KernelSHAP [11] is the most widely used. It transforms the Shapley value estimation into a weighted least squares problem through the Shapley kernel, leading to significantly enhanced estimation efficiency. Based on KernelSHAP, Covert et al. [12] proposed an improved version that can reduce estimation bias, and Jethani et al. [25] developed a computational efficient method called FastSHAP. Additionally, the more complex variants like generalized linear models [26] and kernel regression [27] also belong to regression-based approaches. Compared to sampling-based methods, regression-based methods offer higher computational efficiency and lower estimation variance. However, since most regression methods compute the marginal contribution instead of real contribution of features, they may introduce bias in the estimates.
Gradient-based methods: Gradient-based methods estimate Shapley values by utilizing the gradient of the model’s output with respect to its input features. These methods compute the Shapley value by assessing the change in output resulting from the introduction of each feature increment. DeepSHAP [11] and its variant G-DeepSHAP [28] belong to this category, they are designed to address the computational cost of exact Shapley value estimation. However, DeepSHAP and G-DeepSHAP require access to the gradient of the features, hence are only applicable to differentiable models.
@@
Generative model-based methods: Different from regression-based methods, generative model-based methods estimate the conditional probability of features in Shapley value estimation. For example, the on-manifold Shapley method [29] estimates the Shapley value within the manifold space and employs VAE to compute conditional probabilities across different feature subsets. VAEAC [17], a variant of this approach, applies various masking techniques to the training data to train a model capable of generating any conditional probability distribution. The problem of VAE based generative models is that the amortization error [18] in VAE may bring bias to the Shapley value estimation.
A more comprehensive literature review on Shapley-based feature attribution methods can be found in Ref. [30].
Table I summarizes typical Shapley value feature attribution methods, categorized by their approximation strategy and feature removal technique. It can be seen from Table I that each method has its own advantages, however, they all pose certain limitation in terms of estimation accuracy, which motivated us to undertake the current research.
TABLE I: Summary of typical Shapley value feature attribution methods
Methods Approximation strategy Removal approach Bias-free Pros Cons
ApproShapley[22] Sampling None Simple and general High computation cost
SamplingSHAP[7] Sampling None
SVARM[24] Sampling None
Multilinear extension[23] Sampling Marginal
KernelSHAP[11] Regression Marginal Computationally efficient Assume feature independence
unbiased-KernelSHAP[12] Regression Marginal
FastSHAP[25] Regression Marginal
Generalized additive model[26] Regression Marginal
RKHS-SHAP[27] Regression Marginal
DeepSHAP[11] Gradient Marginal Computationally efficient Model-dependent
G-DeepSHAP[28] Gradient Marginal
On-manifold Shapley[29] Generative model Conditional Flexible Amortization error
VAEAC[17] Generative model Conditional
@@ III Preliminaries
In this section, preliminary knowledge about Shapley value-based explanation framework and energy-based model is presented.
@@ III-B Shapley value-based model explanation
The Shapley value, grounded in a rigorous mathematical framework and underpinned by fairness principles, serves as a powerful tool for feature attribution. By conceptualizing features as players and the model output as the value function, Shapley value offers a principled tool for quantifying the contribution of each feature to the model’s output, thereby enhancing interpretability of complex models.
@@
The output for a specific sample $\mathbf{x}^{t}\in\mathbb{R}^{|D|\times 1}$ in $\mathbf{X}$ can be expressed as the sum of the Shapley values associated with each input feature, obtained by decomposing the model output. Specifically, this decomposition provides the following expression for the output corresponding to $\mathbf{x}^{t}$,
@@
To calculate $\phi_{i}^{t}$, an appropriate contribution function $v(S)$ should be determined. Since $f(\mathbf{x}^{t})$ is the sum of contribution from all input features, we can easily relate $v(\cdot)$ to $f(\cdot)$, that is, $v(D)=f(\mathbf{x}^{t})$ for a specific sample and $\bar{v}(D)=\mathbb{E}[f(\mathbf{x})]$ for global output. Following Ref. [11], the contribution function can be reformed as the expected outcome of $f(\mathbf{x})$ conditioned on the variables in $S$ taking on specific values $\mathbf{x}_{S}^{t}$ as follows.
@@
[…]
@@
where $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ is the conditional density of $\mathbf{x}_{\bar{S}}$ given $\mathbf{x}_{S}^{t}$, $\mathbf{x}_{S}$ and $\mathbf{x}_{S}^{t}$ contains the part of inputs $\mathbf{x}$ and $\mathbf{x}^{t}$ corresponding to features inside $S$, $\mathbf{x}_{\bar{S}}$ and $\mathbf{x}_{\bar{S}}^{t}$ relate to the part of inputs in $\mathbf{x}$ and $\mathbf{x}^{t}$ corresponding to features outside $S$, $S\cup\bar{S}=D$ and $S\cap\bar{S}=\varnothing$.
@@
The Shapley value provides a fair assessment of each variable’s contribution to the model output. However, from Eq. (1) it can be seen that its estimation involves high computation load as there are a total of $2^{|D|-1}$ combinations of $S$. The computational complexity will increase exponentially as the number of input features grows.
III-C Energy-based model
The energy-based models (EBMs) are originated from statistical physics [31]. EBMs are a class of probabilistic model where the probability distribution of data is determined by an energy function. The primary advantage of EBMs is their universal approximation capability [32, 33]. This allows EBMs to model complex, high-dimensional, and multi-modal distributions, capturing intricate dependencies within the data. Unlike traditional parametric models, which rely on restrictive assumptions about the data(e.g., predefined distributions or fixed functional relationships), EBMs do not impose such constraints. As a result, they are well-suited for tasks involving complex or unknown data distributions. Assuming the input features are sampled from an unknown data distribution, i.e., $\mathbf{x}\sim p_{data}(\mathbf{x})$, the goal of energy-based model is to learn an unnormalized density $\exp(-g_{\theta}(\mathbf{x}))$ as a parametric approximation to $p_{data}(\mathbf{x})$.

### v4

@@ Abstract
Shapley value is a widely used tool in explainable artificial intelligence (XAI), as it provides a principled way to attribute contributions of input features to model outputs. However, estimation of Shapley value requires capturing conditional dependencies among all feature combinations, which poses significant challenges in complex data environments. In this article, EmSHAP (Energy-based model for Shapley value estimation), an accurate Shapley value estimation method, is proposed to estimate the expectation of Shapley contribution function under the arbitrary subset of features given the rest. By utilizing the ability of energy-based model (EBM) to model complex distributions, EmSHAP provides an effective solution for estimating the required conditional probabilities. To further improve estimation accuracy, a GRU (Gated Recurrent Unit)-coupled partition function estimation method is introduced. The GRU network captures long-term dependencies with a lightweight parameterization and maps input features into a latent space to mitigate the influence of feature ordering. Additionally, a dynamic masking mechanism is incorporated to further enhance the robustness and accuracy by progressively increasing the masking rate. Theoretical analysis on the error bound as well as application to four case studies verified the higher accuracy and better scalability of EmSHAP in contrast to competitive methods.
@@ I Introduction
Artificial intelligence techniques, particularly deep learning (DL), have been extensively applied to a wide range of complex supervised and unsupervised learning tasks, such as medical diagnosis [1], industrial intelligence [2], and financial modeling [3]. In many scenarios, DL models achieve superior performance compared to conventional methods in terms of accuracy. However, due to their typically black-box nature and complex internal structures, interpreting the results of DL models remains challenging. This opacity hinders the identification of key features influencing model decisions and complicates the understanding of relationships between inputs and outputs. While in tasks where predictive or classification performance is the sole objective, black-box models may be acceptable without requiring interpretability, the situation is markedly different in high-stakes or safety-critical domains such as healthcare, finance, and industrial safety. In these fields, interpretability is not just advantageous but essential for the responsible deployment of artificial intelligence (AI) systems, given the serious consequences that may arise from model errors. The pressing need for transparency in such applications has driven the emergence and growth of Explainable Artificial Intelligence (XAI) as a vital research area [4].
@@
A notable problem for Shapley-based explanation is its high computation load. By definition, the computational complexity of Shapley value increases exponentially with the number of variables. In order to accelerate computation, a series of methods have been developed that can be roughly divided into five categories: sampling-based methods, regression-based methods, gradient-based methods, structure model-based methods, and generative model-based methods. Sampling-based methods utilize Monte Carlo sampling to approximate the true Shapley value. These methods involve Monte Carlo approximation using randomly sampled permutations [11] or sampling based on random permutations of local features [12]. Regression-based methods employ weighted linear regression to accelerate the estimation of Shapley value, with examples including SHAP (SHapley Additive exPlanation) [13] and KernelSHAP [14]. Gradient-based methods estimate Shapley values by exploiting the gradient of the model output with respect to input features. Typical approaches include DeepSHAP [13] and its variant G-DeepSHAP [15]. Structure model-based methods leverage specific model structure constraints to approximate the Shapley value. Examples of such methods include tree Shapley [16] and graph Shapley [17]. More recently, generative model-based methods have been developed for Shapley value estimation. A representative example is the variational autoencoder with arbitrary conditioning (VAEAC) [18], which uses a single model for all feature combinations.
@@
Despite the advancements made in the calculation of Shapley value, several significant issues still remain. For example, sampling-based methods require a large quantity of samplings to achieve sufficient accuracy, and it may suffer from high variance during sampling. Regression-based methods can improve the estimation efficiency of Shapley value, however, the inherent assumption of independence among variables may negatively affect its accuracy. Whilst gradient-based and structure model-based approaches demonstrate improved efficiency, their application scenarios are limited to specific models. Another notable research gap is that there still lacks a unified theoretical framework on the error bounds of different Shapley value-based approaches, making it difficult to quantitatively evaluate and compare their accuracy.
@@
In order to increase estimation accuracy and provide theoretical guidance on the error bound, this article presents the EmSHAP (Energy-based model SHapley value estimation). Similar to Ref. [18], EmSHAP introduces a generative model during the estimation of Shapley contribution function. It relies on the energy-based model(EBM) to obtain the expectation of Shapley contribution function of the deep learning model under some arbitrary subset of features given the rest. The use of EBM is motivated by its ability to model complex, high-dimensional distributions without explicit normalization, which provides a flexible and efficient framework for estimating the conditional distributions required in Shapley value computation. Unlike original EBM, the EBM in EmSHAP is decomposed into an energy network to approximate the unnormalized conditional density and a GRU network for approximation of the proposal distribution in the partition function. The GRU network has a simple structure and is capable of capturing long-term dependencies, and the adopted dynamic masking scheme improves the flexibility. The contributions of the paper can be summarized as: i) Methodological refinements on the EBM framework by introducing the GRU network and the dynamic masking scheme, resulting in more accurate Shapley value estimation; ii) Theoretical analysis of error bounds for several state-of-the-art methods are performed with the purpose of providing a quantitative measure to evaluate and compare the accuracy of different methods; iii) Extensive application experiments are conducted to verify the accuracy and scalability of different Shapley value estimation methods, so that practitioners can refer to when choosing appropriate Shapley value estimation methods for specific applications.
II Related work
@@
Shapley value is a concept originated from cooperative game theory that provides a fair and reasonable method for quantifying feature attribution contributions. As a result, they have garnered significant attentions in the field of XAI. However, the calculation of exact Shapley values is an NP-hard problem [19]. To address this computational complexity, researchers have developed various approximation methods. This section provides a brief review on the different kinds of approximation methods in the literature, which can be summarized into sampling-based, regression-based, gradient-based, structure model-based and generative model-based methods.
@@
Sampling-based methods: Sampling-based Shapley value estimation is among the earliest approaches studied and remains the most widely adopted in practice. The core principle is to randomly sample subsets of features and compute the average contribution of the features to the model output. The most commonly used sampling-based methods are Monte Carlo sampling [20, 21], and Monte Carlo sampling with random arrangements [22, 7]. In addition to standard Monte Carlo sampling, recent studies have explored alternative strategies to improve estimation accuracy, such as Owen-based sampling [23], Quasi-Monte Carlo sampling [24], permutation sampling [25], and stratified sampling [26, 27]. Most of these approaches are model-agnostic and do not require explicit feature removal, making them attractive for practical applications. However, sampling-based methods may suffer from high variance during the sampling procedures.
@@
Regression-based methods: Regression-based methods estimate Shapley values by using regression models to approximate the marginal contribution of each feature to the output. These methods approximate the original black-box model by constructing a simpler surrogate model. Among these methods, KernelSHAP [13] is the most widely used. It transforms the Shapley value estimation into a weighted least squares problem through the Shapley kernel, leading to significantly enhanced estimation efficiency. Subsequent works further extended this line of research. For example, Covert et al. [14] proposed an improved version of KernelSHAP to reduce the estimation bias, Simon and Vincent [28] introduced SGD-Shapley, which uses a stochastic gradient descent scheme to iteratively solve the weighted least squares problem, while Jethani et al. [29] developed FastSHAP, a more computationally efficient variant ST-SHAP [30] is developed to enhance stability by employing layer-wise neighbor selection. Additionally, the more complex variants like generalized linear models [31] and kernel regression [32] also belong to regression-based approaches. Compared to sampling-based methods, regression-based methods offer higher computational efficiency and lower estimation variance. However, since most regression methods compute the marginal contribution instead of the real contribution of features, they may introduce bias in the estimates.
Gradient-based methods: Gradient-based methods estimate Shapley values by exploiting the gradient of the model output with respect to input features. The basic idea is to approximate the contribution of each feature by evaluating how incremental changes in the feature influence the output. Typical approaches include DeepSHAP [13] and its variant G-DeepSHAP [15], both designed to alleviate the computational cost of exact Shapley value estimation. In addition, Ancona et al. [33] proposed a polynomial-time approximation for Shapley values in deep neural networks based on uncertainty propagation, and Wang et al. [34] introduced Shapley Explanation Networks (SHAPNETs) that treat Shapley values as latent representations to enable hierarchical explanations and regularization. Despite their efficiency, gradient-based methods require access to model gradients and are therefore limited to differentiable architectures.
Structure model-based methods: Structure model-based methods estimate Shapley values by leveraging the inherent structure of specific models to simplify computation. These methods exploit the constraints and decomposition properties of structured models, such as decision trees and graphs, to efficiently calculate exact or approximate Shapley values. Tree Shapley [16], TreeSHAP [35], and graph Shapley [17] fall into this category. They utilize the model structure to significantly reduce computational complexity. However, structure model-based methods are restricted to certain model classes so that their application is limited.
@@
Generative model-based methods: Different from other methods, generative model-based methods estimate the conditional probability of features in Shapley value estimation. For example, the on-manifold Shapley method [36] estimates the Shapley value within the manifold space and employs VAE to compute conditional probabilities across different feature subsets. VAEAC [18], a variant of this approach, applies various masking techniques to the training data to train a model capable of generating any conditional probability distribution. The problem of VAE-based generative models is that the amortization error [37] in VAE may bring bias to the Shapley value estimation.
In addition to the above categories, recent studies have extended Shapley values to quantify feature interactions rather than individual feature importance. Techniques like SHAP-IQ [38], TreeSHAP-IQ [39], KernelSHAP-IQ [40], and SVARM-IQ [41], which provide effective frameworks to identify and interpret higher-order interactions among features. These methods are particularly useful in settings where feature dependencies play a central role.
The aforementioned methods represent the mainstream approaches for Shapley value estimation. For a broader synthesis, readers are referred to Ref. [42].
@@ III Preliminaries
In this section, preliminary knowledge about Shapley value-based explanation framework is presented.
@@ III-B Shapley value-based model explanation
The Shapley value, grounded in a rigorous mathematical framework and underpinned by fairness principles, serves as a powerful tool for feature attribution. By conceptualizing features as players and the model output as the value function, Shapley value offers a principled tool for quantifying the contribution of each feature to the model’s output, thereby enhancing the interpretability of complex models.
@@
The output for a specific sample $\mathbf{x}^{t}\in\mathbb{R}^{|D|}$ in $\mathbf{X}$ can be expressed as the sum of the Shapley values associated with each input feature, obtained by decomposing the model output. Specifically, this decomposition provides the following expression for the output corresponding to $\mathbf{x}^{t}$.
@@
To calculate $\phi_{i}^{t}$, an appropriate contribution function $v(S)$ should be determined. Since $f(\mathbf{x}^{t})$ is the sum of contribution from all input features, we can easily relate $v(\cdot)$ to $f(\cdot)$, that is, $v(D)=f(\mathbf{x}^{t})$ for a specific sample and $\bar{v}(D)=\mathbb{E}[f(\mathbf{x})]$ for global output. Following Ref. [13], the contribution function can be reformed as the expected outcome of $f(\mathbf{x})$ conditioned on the features in $S$ taking on specific values $\mathbf{x}_{S}^{t}$ as follows.
@@
$=\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d{\mathbf{x}_{\bar{S}}},$
@@
$\bar{S}$ is the complement of subset $S$ in the feature set $D$ (i.e., $\S\bar{S}=D\backslash$ S), $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ is the conditional density of $\mathbf{x}_{\bar{S}}$ given $\mathbf{x}_{S}^{t}$, $\mathbf{x}_{S}$ and $\mathbf{x}_{S}^{t}$ contains the part of inputs $\mathbf{x}$ and $\mathbf{x}^{t}$ corresponding to features inside $S$, $\mathbf{x}_{\bar{S}}$ and $\mathbf{x}_{\bar{S}}^{t}$ relate to the part of inputs in $\mathbf{x}$ and $\mathbf{x}^{t}$ corresponding to features outside $S$.
@@
The Shapley value provides a fair assessment of each feature’s contribution to the model output. However, from Eq. (1) it can be seen that its estimation involves high computation load as there are a total of $2^{|D|-1}$ combinations of $S$. The computational complexity will increase exponentially as the number of input features grows.
IV Methodology
By definition of Shapley value in Eq.(1), it can be seen that the most important step is to estimate the contribution function $v(S)$, which can be reformulated as the conditional expectation in Eq.(3). The conditional expectation can be empirically approximated by techniques like Monte Carlo integration provided the conditional probability $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ is known.
$v(S)$ $=\mathbb{E}[f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S})|\mathbf{x}_{S}=\mathbf{x}_{S}^{t}]=\int f(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}^{t})p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})d\mathbf{x}_{\bar{S}}$ (4)
where $\mathbf{x}_{\bar{S}}^{(k)}\sim p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S}=\mathbf{x}_{S}^{t})$ for k=1,2,,Kk=1,2,...,K, and $K$ is the number of Monte Carlo sampling. It is worth noting that estimation of the conditional density $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ can be achieved by techniques like KDE. However, KDE is computationally challenging for high-dimensional inputs. More importantly, since there are $2^{|D|-1}$ possible subsets of features, KDE would require estimating an exponential number of conditional densities, rendering it infeasible for practical applications.
IV-A Network structure
To address the computation challenge of conditional density estimation, EmSHAP is designed to capture the complex dependencies among features and produce accurate conditional estimates, which are then used to calculate the Shapley values. The detailed network structure of EmSHAP is shown in Fig. 1.
Figure 1: Structure of EmSHAP for estimating Shapley value.
The network structure in Fig. 1 mainly consists of three components, namely, the data masking mechanism for generating different combinations of conditional densities, the energy network for estimating of energy function/unnormalized density function and the GRU network for estimation of proposal distribution. The proposal distribution serves to estimate the partition function, which provides the normalization for the energy function. In Fig. 1, data batches of input features are masked and fed into the energy network to estimate the energy function $g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S})$ and the unnormalized probability $\text{exp}(-g_{\theta}(\mathbf{x}_{\bar{S}},\mathbf{x}_{S}))$, the unnormalized probability approximates the conditional probability $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ after divided by the partition function. The energy network is designed as an unnormalized Multilayer Perceptron (MLP) with skip connections [43]. The partition function, on the other hand, needs to be approximated by Monte Carlo integration from the proposal distribution $q(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$. The proposal distribution can be estimated using an autoregressive approach [44], however, its estimation accuracy will be negatively affected by feature orderings [^1] Feature ordering refers to the arrangement of input features in a specific sequence when fed into a neural network.. Instead, a GRU network with a dynamic masking scheme is used, which outputs parameters related to the estimated proposal distribution and a context vector $\bm{\gamma}$. The context vector is used to connect the energy network and the GRU network. With the conditional probability $p(\mathbf{x}_{\bar{S}}|\mathbf{x}_{S})$ obtained, the contribution function can be estimated from Eq.(4), and Shapley value can be subsequently obtained from Eq.(1). This network structure is under the framework of energy-based model, which can universally approximate any probability density according to Ref. [45]. Hence the proposed model structure inherits the good approximation ability of EBM. It should be noted that for estimating the proposal distribution, other RNN variants, such as Long Short-Term Memory (LSTM), BiGRU, and Attention-based GRU (Att-GRU), can also be applied. However, the more complex gating structure of LSTM often leads to higher computational costs and an increased risk of overfitting. On the other hand, advanced variants such as BiGRU and attention-augmented GRU (Att-GRU) require substantially more time and memory during training and inference, while the observed performance improvement may be marginal. This is because the introduction of bidirectional or attention mechanisms may introduce redundant information that does not substantially enhance feature attribution, and the added complexity can exacerbate the risk of overfitting. Hence, the GRU is selected over other RNN variants, as it represents a more balanced choice that offers a favorable trade-off between performance and computational efficiency. This conclusion is further supported by the comparative studies in Section VI-E. By capturing the intrinsic information associated with the sequential masking of variables, the GRU network ensures that the generated proposal distribution is more accurate, thereby enabling the model to better approximate complex conditional probability densities. The details of the network structure will be explained in Subsections IV-B and IV-C.
IV-B Energy network for estimation of unnormalized conditional probability density
For estimation of Shapley values, EmSHAP relies on the energy-based model. EBMs are a class of probabilistic models with the probability of a configuration $\mathbf{x}$ defined through an energy function $g_{\theta}(\mathbf{x})$ [46]. EBMs have strong universal approximation capability [47, 48], enabling them to capture high-dimensional and multi-modal distributions without restrictive assumptions on the data. Formally, the data distribution can be approximated as follows.